OpenStudy (anonymous):

How can I solve this? Estimate the indicated derivative by any method. (Round your answer to three decimal places.) g(t)=6/t^5; estimate g'(−1)

OpenStudy (mathstudent55):

\(g(t) = 6/t^5\) \(g(t) = 6t^{-5}\) \(g'(t) = -5 \times 6t^{-6}\) \(g'(t) = -30/t^{6}\) Now plug in -1 for t.

OpenStudy (anonymous):

I got 30 but that's not right

OpenStudy (mathstudent55):

What is \((-1)^6\)?